A school club is planning a bus trip to the wildlife park. Rental fee for a bus which will hold up to 45 people is 600 centos (units of money) and admission tickets cost 30 centos each. If the cost of the trip, including bus and admission ticket, is set at 50 centos per person, what is the minimum number of people who must participate to ensure that these costs are covered.
A. 12
B. 20
C. 24
D. 30
E. 45

Answer by Theo(13342)

the cost of the bus is 600.
the cost per student is 30.
what is collected from each student is 50.
let x = number of students.
revenue collected from each student is equal to 50x.
cost of the trip is equal to 600 + 30x.
to break even the cost must be equal to the revenue.
you get:
50x = 600 + 30x
subtract 30x from both sides of this equation to get:
20x = 600
divide both sides of this equation by 20 to get:
x = 30
the breakeven point is 30 students.
more than that and a profit is made.
less than that and a loss is incurred.
you can see that from the following table:
number of students    revenue              cost               net
       20             20*50=1000          600+20*30=1200      loss of 200
       30             30*50=1500          600+30*30=1500      break even
       40             40*50=2000          600+40*30=1800      profit of 200